OpenStudy (anonymous):

5 (x • y) = (5 • x) y is an example of which algebraic property?

OpenStudy (anonymous):

Pick out of these. Commutative Property of Multiplication: \(5\times3\times2 = 2 \times 5 \times 3\) Associative Property of Multiplication: \(5(3 \times 2) = (5 \times 3)2\) Zero Property of Multiplication: \(0 \times 2 = 0\) Identity Property of Multiplication: \(1 \times 2 = 2\)

OpenStudy (anonymous):

Commutative Property of Multiplication

OpenStudy (anonymous):

No. Think again.

OpenStudy (anonymous):

Associative Property of Multiplication

OpenStudy (anonymous):

Yup :)